# Static and Dynamic BART for Rank-Order Data

### v3

@@ 2.3 Dynamic Model: AutoRegressive Rank-Order BART
$p(\mathbf{z}_{t}|\tau_{1:t})\propto\mathbb{I}(\mathbf{z}_{t}\in A_{t})\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{k}\in\{1,\ldots,K\}^{N}}\mathcal{N}_{N}(\mathbf{z}_{t}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N})q_{\mathbf{k},t}(A_{1:t-1}),$ (7)
@@
$q_{\mathbf{k},t}(A_{1:t-1})$ $=\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{m}\in\{1,\ldots,K\}^{N}}q_{\mathbf{m},t-1}(A_{1:t-2})\int_{\mathbf{C}_{\mathbf{k},t-1}}\mathcal{N}_{N}(\mathbf{z}_{t-1}|\bm{\tilde{\mu}}_{\mathbf{m}},I_{N})\>\mathrm{d}\mathbf{z}_{t-1},$
@@
$p(\mathbf{z}_{t+1}|\tau_{1:t})\propto\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{k}\in\{1,\ldots,K\}^{N}}\mathcal{N}_{N}(\mathbf{z}_{t+1}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N})q_{\mathbf{k},t+1}(A_{1:t}).$ (8)
@@
$p(\mathbf{z}_{t}|\tau_{1:t})$ $\propto\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{k}\in\{1,\ldots,K\}^{N}}w^{F}_{\mathbf{k},t}u^{F}_{\mathbf{k},t}(\mathbf{z}_{t}),$ (9)
@@
$w^{F}_{\mathbf{k},t}$ $=\frac{q_{\mathbf{k},t}(A_{1:t-1})n_{\mathbf{k},t}}{\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{m}\in\{1,\ldots,K\}^{N}}q_{\mathbf{m},t}(A_{1:t-1})n_{\mathbf{m},t}}.$
@@
$p(\mathbf{z}_{t}|\tau_{1:T})$ $\propto\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{k}\in\{1,\ldots,K\}^{N}}w^{S}_{\mathbf{k},t}u^{S}_{\mathbf{k},t}(\mathbf{z}_{t}),$ (10)
@@
$w^{S}_{\mathbf{k},t}$ $=\frac{q_{\mathbf{k},t}(A_{1:t-1})m_{\mathbf{k},t}}{\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\ell}\in\{1,\ldots,K\}^{N}}q_{\bm{\ell},t}(A_{1:t-1})m_{\bm{\ell},t}},\qquad\mathbf{k}\in\{1,\ldots,K\}^{N}.$

### v4

@@ 2.3 Dynamic Model: AutoRegressive Rank-Order BART
$p(\mathbf{z}_{t}|\tau_{1:t})\propto\mathbb{I}(\mathbf{z}_{t}\in A_{t})\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{k}\in\{1,\ldots,K\}^{N}}\mathcal{N}_{N}(\mathbf{z}_{t}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N})q_{\mathbf{k},t}(A_{1:t-1}),$ (7)
@@
$q_{\mathbf{k},t}(A_{1:t-1})$ $=\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{m}\in\{1,\ldots,K\}^{N}}q_{\mathbf{m},t-1}(A_{1:t-2})\int_{\mathbf{C}_{\mathbf{k},t-1}}\mathcal{N}_{N}(\mathbf{z}_{t-1}|\bm{\tilde{\mu}}_{\mathbf{m}},I_{N})\>\mathrm{d}\mathbf{z}_{t-1},$
@@
$p(\mathbf{z}_{t+1}|\tau_{1:t})\propto\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{k}\in\{1,\ldots,K\}^{N}}\mathcal{N}_{N}(\mathbf{z}_{t+1}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N})q_{\mathbf{k},t+1}(A_{1:t}).$ (8)
@@
$p(\mathbf{z}_{t}|\tau_{1:t})$ $\propto\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{k}\in\{1,\ldots,K\}^{N}}w^{F}_{\mathbf{k},t}u^{F}_{\mathbf{k},t}(\mathbf{z}_{t}),$ (9)
@@
$w^{F}_{\mathbf{k},t}$ $=\frac{q_{\mathbf{k},t}(A_{1:t-1})n_{\mathbf{k},t}}{\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{m}\in\{1,\ldots,K\}^{N}}q_{\mathbf{m},t}(A_{1:t-1})n_{\mathbf{m},t}}.$
@@
$p(\mathbf{z}_{t}|\tau_{1:T})$ $\propto\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{k}\in\{1,\ldots,K\}^{N}}w^{S}_{\mathbf{k},t}u^{S}_{\mathbf{k},t}(\mathbf{z}_{t}),$ (10)
@@
$w^{S}_{\mathbf{k},t}$ $=\frac{q_{\mathbf{k},t}(A_{1:t-1})m_{\mathbf{k},t}}{\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\ell}\in\{1,\ldots,K\}^{N}}q_{\bm{\ell},t}(A_{1:t-1})m_{\bm{\ell},t}},\qquad\mathbf{k}\in\{1,\ldots,K\}^{N}.$